A sample solution containing sulfate ions (SO42−) was analyzed and was found to contain 5 mol of sulfate ions (SO42−).
Calculate the equivalent of sulfate ions (SO42−) present in the solution.
1
Expert's answer
2018-11-29T12:17:37-0500

Answer on Question #83431, Chemistry/ Organic Chemistry

A sample solution containing sulfate ions (SO42−) was analyzed and was found to contain 5 mol of sulfate ions (SO42−).

Calculate the equivalent of sulfate ions (SO42−) present in the solution.

Solution

Equivalent weight of sulphate = 96.07/2= 48.04 g/equivalent

5 mol will have a weight of = 5*96.07 = 480.35 g

The equivalent of sulfate ions = 480.35/48.04 = 10.00 eq

Answer: 10.00 eq